40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 13 03:22:53 2011, kiwamu, Update, LSC, measurement of sensing matrix : just began 
    Reply  Fri Oct 14 04:35:45 2011, kiwamu, Update, LSC, locking tonight Screen_shot_2011-10-14_at_4.24.01.png
Message ID: 5659     Entry time: Thu Oct 13 03:22:53 2011     Reply to this: 5665
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: measurement of sensing matrix : just began 

- status update on LSC activity :

The measurement of the LSC sensing matrix has begun. But no useful results yet.

 

 The measurement script (#4850) ran pretty well after I did some modifications to adopt the script to the latest LSC model.

However the SNR weren't so great particularly in REFL33 in the PRMI configuration.

So I will tune the amplitude of excitations and integration times tomorrow.

Currently the excitation is at 238.1 Hz, where no disturbing structures are found in the spectra.

ELOG V3.1.3-